Circular No. 6126

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

RE J1255+266 M. G. Watson, University of Leicester; R. G. McMahon, Institute

of Astronomy, Cambridge; and M. J. Page, MSSL, University College, London, report: “We looked for an optical counterpart to the extremeultraviolet (EUV) transient RE J1255+266 (IAUC 6085) on Jan. 12 with the William Herschel 4.2-m telescope (+ ISIS spectrograph). We propose as the counterpart a blue star of mag V =18.5, lying 72′′ from the reported position of RE J1255+266, which is outside the quoted error circle (ibid.). Nevertheless, the unusual characteristics of this star strongly suggest that it is the correct counterpart. It is the bluest optical object visible on the Palomar Sky Survey within a radius of 180′′. From the APM measurement of the POSS, the counterpart is located at α = 12^{h}55^{m}10^{s}.7, δ = +26^{o}42′28′′ (equinox 2000.0, epoch 1955.3, uncertainty ± 1′′). The optical spectrum shows a blue continuum together with broad Balmer hydrogen absorption lines, characteristic of a DA white dwarf. In addition, very strong, narrow Balmer emission lines are seen. If the identification is correct, this spectrum suggests that RE J1255+266 is a binary system with a low-mass companion heated by the white dwarf. The EUV outburst of this object might be associated with a brief episode of accretion onto the white dwarf if the system is close to being semi-detached, suggesting a short orbital period of < 4 hr. Spectroscopic and photometric observations are urged.”

NOVA CASSIOPEIAE 1993 D. K. Lynch, R. W. Russell, R. C. Kellogg, and A. L. Mazuk, The

Aerospace Corporation; and M. S. Hanner, Jet Propulsion Laboratory, report on their spectroscopy (range 3–14 μm, resolution 30–130) from the Infrared Telescope Facility at Mauna Kea: “On 1994 Dec. 14.3 UT, the nova showed a strong infrared excess, with infrared magnitudes L = 6.6 ± 0.2, M = 5.5 ± 0.07, and N = 1.8 ± 0.15. The continuum between 8 and 13 μm falls off more steeply than a Planck function. The spectrum displayed two prominent emission features, one at 3.4 μm (∼ 0.25 μm wide) and the other at 11.3 μm (∼ 0.4 μm wide). There was also evidence for an unresolved emission line at 8.8 μm. If the 3.4- and 11.3-μm emissions are due to those that cause unidentified infrared features, the 8.6-μm feature is noticeably absent.”

1995 January 16 Daniel W. E. Green